# Dynamic Neural Curiosity Enhances Learning Flexibility for Autonomous Goal Discovery

### v1

@@ III-E Curiosity
When the robot discovers a new goal, a forward and an inverse model are created. The forward model output the first error in order to compute the learning progress with dynamic neural fields (Fig. 8). During exploitation, the robot performs a motion and observes the outcomes in the environment. If it results to no changes (the object did not move), then the error remains the same. In case of any changes, the error controller computes the error with the forward model’s prediction. Here, the model follows a precise timing for the error computation where the observation takes place ahead of the end of the motion. In practice, this mean the robot pushes the object, computes the error while going back to a resting position and finally signals that the motion is done. The curiosity mechanism is then divided into two parts : the error and learning progress processes. The former focuses on characterizing the error of a goal as neural activity while the latter computes the error’s level dynamic to deliver a learning progress. We will first focus on the error architecture before presenting the learning progress structure.
Figure 8: Architecture of the error component. The process formalizes the error coming from the forward model of a specific goal as neural activation.
The compute the error NF gathers error levels from all the goals discovered. When the robot pursues the learning of a skill, the error level from goal error MT is combined with the filter MT in order to only affect the current goal. Indeed, this last field regroups each error and inhibits the current goal whereas the goal error MT provides a single peak (i.e., the goal error’s activation). It is worth to mention that the activation function ($\sigma$) used for this field, local goal NF, local error NF and updated error NF is a ReLU function. The time node acts as a trigger when a new error value is computed. Once a new error is delivered by the error controller, this brief activation supports the update of the memory trace as well as boosting the resting level of other neural fields. The errors MT is there to keep in memory each goal value and only updates them once the end of the action node (EOA) is active. The last main fields are the local goal NF and updated error NF. They have a resting level of $-2$, meaning that they yield peaks of activation only when they receive the boost from the time node. As mentioned earlier, the error is computed before the robot reaches its resting pose and activates the EOA node. This implies that when the time node is active, the local error NF contains the error’s goal before the update whereas the updated error NF delivers the newly calculated value. These last fields prepare the future neural computation of the various learning progresses.
Figure 9: Neural flow of the learning progress. The module processes neural dynamics coming from the error component to generate a learning progress.
@@
The learning progress mechanism is straightforward (Fig. 9). It receives inputs from updated error NF and local goal NF that are projected into zero dimensional nodes with resting level of $-1.025$. The time node is the same seen in the error architecture. With this pattern of incoming activation, the rise and low nodes are responsible for detecting if the updated error value is respectively superior or inferior to the previous level of activation (local goal NF). The filter NF follows the same purpose as the one described in the error architecture, providing the error level of each goal except the current one. Then, the compute LP memory trace is where the learning progress of a goal is computed. In the event where the low node is active, the supra-threshold activation from variation NF will project into compute LP and build up an activation. When the rise node is active, it indicates that the error level is rising. In that case, there will be no activation within variation NF and compute LP will decay the learning progress of the current goal. If an activation was present within variation NF, the activation will rise depending on the value of ($\tau_{+}$). Finally, there is a case where the values of the error before and after update are too close to each other, which signify that the progress is not significant. Then, the rise and low nodes will not activate. The specific resting level value of these nodes support the fine tuning of this specific event. Here, there will be no activation if the error does not vary beyond a threshold of 0.025.
@@ IV Experimental setup
The purpose of the experiments is to evaluate the learning behavior of the architecture under different parameters and with different objects. More specifically, we want to determine the role of habituation, persistence and error inhibition during goal discovery and learning. We will also analyse the tonic and phasic activation within the LC. Then, we will assess how the Locus Coeruleus model gathering bottom-up attention and curiosity lead to an open-ended learning approach.
@@
The scene simulates a Franka Emika collaborative robot in Gazebo (Fig. 10). The dynamic neural fields are simulated with the software Cedar [70] and the middleware ROS [71] insures the communication. The objects used for the setup are a red ball, a blue cube and a yellow cylinder. We specifically choose these objects in regard to their difficulty to be manipulated. Since the end-effector orientation is fixed with the gripper closed, it is relatively easy to be able to push the cube in different directions. For the cylinder, the last pose generated has to be more precise to avoid seeing the object slip away during a motion. The ball is the most difficult object to learn to manipulate since a slight error in the end effector position can push the ball in various direction. The experiment begins when an object is present in the camera visual field. The object is respawned at the center of the table if it is out of reach from the robot. It means that the objects can be moved and rotated from various location on the table as long as it is not out of reach. Each neural field is bounded between 0 and 100 along its respective dimension. The gripper orientation is fixed as well as the z dimension while performing a motion. During the goal discovery stage, the end-effector position of the robot is recorded from the resting state until reaching the goal pose (III). The number of recorded poses remains almost the same, depending on the motion ($\approx$ 30 poses) and they are used to generate a DMP if the motion led to a new stimuli.
@@
The visual perception is processed by a virtual camera with a resolution of 1280$\times$720. This camera is calibrated to the robot work space in order to transform the position of an object from the camera point of view to the robot base. A tracker identifies the object by color and returns its position to the position object NF in the action module. At the same time, the color tracker returns a continuous input to several neural fields (perception and habituation modules). As mentioned earlier, a goal is formalized within a neural field by collecting the motion direction and color of the object along the horizontal and vertical dimensions respectively.
@@
The learning of a skill is formalized by a forward and inverse model through multi-layer perceptrons. The forward model has one input layer (4D), one hidden layer (6D) and one output layer (1D). The inputs are the current position of the object (x,y) and the motor command (final pose of the end-effector $x,y$). Then, the output of this neural network is the motion (goal angle) of the object. The inverse model has one layer input (3D) composed of the current object position (x,y) and the goal angle of the object. After the hidden layer (4D), the output (2D) is the motor command (end-effector position on x and y). Each DMP is associated with its own forward and inverse model. The purpose of the inverse model associated with a DMP is to learn to generate the correct end-effector final pose for a given goal and object position. More details of the learning parameters can be found in Appendix -F.
@@ V-A Habituation
In this section, we evaluate the effect of habituation during exploration (see Fig. 11). To do so, we performed ten explorations for each of the three objects (cube, cylinder and ball) by varying the $\tau_{+}$ (Eq. 4) parameter of the visual memory trace.
Figure 11: (Top) Average number of goals discovered for 10 experiments on each object with fast and slow habituation. The values for fast and slow habituation ($\tau_{+}$) are 2 and 4 seconds.
As expected, a slow habituation ($\tau_{+}=4$ seconds) allows the robot to spend more time on exploration, leading to an increase in the number of goals discovered (see Fig. 11a). We can notice a difference among the objects with the ratio :
@@
Figure 12: Neural activation representing the evolution of the learning progress in time with standard deviation across 20 experiments for the baseline experiment. The figure depicts the average learning progress for the cube (left), the cylinder (middle) and the ball (right).
where $S_{avg}$ and $F_{avg}$ are the average number of goals discovered for a slow and fast habituation. The ratio for the cube, cylinder and the ball are respectively 1.53, 2.04 and 1.69. This suggest that given enough habituation time, the robot can double the number of goals discovered with the cylinder. This can be explained by the orientation of the gripper and the inhibition of return. Since the gripper’s orientation is fixed, only a limited number of actions can trigger the discovery of a new goal. The movement is guided by the inhibition of return around the object’s location, so this inhibition has to be precise and thus requires more time to generate a meaningful motion. In the following section, we will analyse the learning trajectory of different goals with curiosity.
V-B Goal learning progress
@@
Skill sequence Cube Cylinder Ball Cube Cylinder Ball Cube Cylinder Ball
@@
Table I: Mann-Whitney U test between 2 consecutive skills for each object, under each setting and for all experiments.
@@ V-B1 Time-course of learning progress between objects
The evolution of the learning progress is presented in Fig. 12. We run the experiment twenty times for each object and, to compare the learning between objects and parameters, we fixed the same four goals. These goals are : push up, push to the left, push to the right, push down. The weights initialization of the forward models determine the first error during goal discovery, which then influence the future goal selection for learning. We consider this experiment as the baseline, with a persistence value (transient action MT node) at $4500ms$ and $100ms$ for respectively $\tau_{+}$ and $\tau_{-}$. The error inhibition (attempts MT projected on LC:phasic) is settled at $100ms$ for $\tau_{+}$ and $\tau_{-}$, which means there is no error inhibition staying in memory between the selection of a goal to pursue. The learning progress (LP MT in learning progress module) is fixed at $2000ms$ for $\tau_{-}$ and $\tau_{+}$. Regarding the LP, we arbitrarily choose these values because a high $\tau_{-}$ would signify that the goal error takes more time to decay even if the error coming from the forward model is already steady at a low level. We performed a Kruskal-Wallis test to evaluate the significance of the skills ordering for each object. For the cube, we took the distribution from $t=[0,1800]s$; For the cylinder, we sample $t=[0,2250]s$; For the ball, the distribution range is $t=[0,6000]s$. The differences of distribution between the four goals are significant with $H(4)$, $p<0.01$ among the three objects. Even if the sequence of the four goals is significant among all the runs, we still perform a Mann-Whitney U test (see Tab. I) between two successive goals (push-up $\rightarrow$ push-left, push-left $\rightarrow$ push-right, push-right $\rightarrow$ push-down) to verify a possible overlap between them in certain runs. The statistical test for the cylinder is almost not significant, which indicates a different ordering between $up$ and $left$ among several runs. We can conclude that the goals follow the same order of learning for the baseline experiment. However, these skills do not reflect a developmental trajectory since they can be learned independently. Moreover, the learning begins with the goal that has the highest error, which depends on the synaptic weights initialization of the forward model. The difficulty of a goal to be learned can be evaluated by the time spent until the learning progress activation comes below $0.01$. For the ball, it is considerably more difficult to learn the set of goals than the cylinder and the cube ($\approx$ 6000s). This is essentially due to the precision required to generate a $goal$ (Appendix -B).
@@ V-B2 Winner inhibits similar goals
Figure 13: (Left) Three goal errors in errors MT belonging to the same object where two of them are similar. The highest peak is selected as current goal for learning when there are no learning progress. (Right) Effect of the inhibition of the current goal on its neighborhood inside filter MT.
During the exploration with the habituation phase, it is possible for the robot to discover similar goals. For example, the robot can push a ball up and push a ball up but slightly shifted by a few degrees. In that case, only the goal with the highest error coming from the forward model will be learned in the future. This comes from the neural dynamics of the filter MT (Error module), and more particularly from the projection of the current goal (see Fig. 13). Indeed, the filter MT represents the various goal errors discovered except the goal being currently learned via the inhibition of current goal NF. This localized inhibition not only affects the goal peak, but also partially or completely alters any peaks of activation at proximity. This effect depends on the standard deviation ($\sigma$) of the inhibition. A nearby goal will be affected by having its peak level lowered, then updated within the compute error NF. This essentially indicates that each time the error is computed for the current goal, any adjacent goals will see a fall on their activation level. A higher standard deviation for the peak activation or a wider inhibition value from current goal NF means that the goals must significantly differ from each other. Indeed, it can be seen as a way to first focus on acquiring a diverse set of simple skills before specializing. The consequences of this effect lead to developmental questions and should be investigated more deeply.
@@ V-C Influence of inhibition and persistence
After the learning of goals under the baseline parameters, we are adjusting the error inhibition dynamics that will be projected to the LC (via attempts MT). We settle the time parameters at $\tau_{+}=2000ms$ and $\tau_{-}=1500ms$ and run the experiment twenty times for each object (see Fig. 14, top row). We arbitrarily chose these parameters so the inhibition takes more time to build up and decay. For a goal where the error is difficult to decrease, it will deliver a mild inhibition that would avoid the goal to be selected again in a near future. We repeated a Kruskal-Wallis test for the cube, cylinder and ball for the time intervals $t=[0,2100]s$, $t=[0,2500]s$ and $t=[0,4250]s$ respectively. The distribution of the four skills across the 20 experiments was conclusive with $H(4),p<0.01$. A Mann-Whitney test was performed to investigate if the goals are following the same order consecutively (see Table I). We observed that the test is not significant for the cube, which signifies that there is no clear ordering between $pushing-up$ and $pushing-left$. Indeed, the $pushing-up$ goal being difficult to learn, if the learning progress did not rise after a run of learning, the inhibition will prevent the selection of this goal in the near future. The statistical test result was significant for the cylinder and the ball but with a higher value than the baseline experiment. This demonstrates the presence of several runs for both objects where the goals do not follow the mainstream sequence pattern. If we compare the inhibition settings with baseline, we can see a rise in the learning time for the cube and cylinder. However, this learning time is shorter for the ball and the goals $push-right$ and $push-down$ are less subject to fluctuation. For the ball baseline experiment, $push-right$ and $push-down$ take respectively $5000s$ and $3900s$ against $3000s$ and $2000s$ under the inhibition parameters. The baseline experiment only relies on the learning progress of the goals to continue the exploitation, which provokes only minimal learning fluctuation in case of easy objects to learn (cube and cylinder). For a difficult object, the system has to continue the learning of hard goals that exhibits a certain learning progress with no possibility to select a different skill to exploit. Under the inhibition parameters, the model introduces more flexibility in the goal exploitation since the learning progress is not the only factor of selection. Essentially, this suggests that the robot does not finish learning a goal entirely before learning a new one. This switch between goals indicates an advantage on the ball whereas it introduces more fluctuation for the cube. Apart from a certain learning time extension ($\approx 250s$), the cylinder does not seem to be affected by the inhibition parameters.
@@
We introduce the persistence settings by modifying the transient action MT node ($\tau_{+}=6000ms$, $\tau_{-}=100ms$). These settings are extending the time of exploitation before the LC module updates its decision regarding the goal to learn. This parameter is analog to the exploration time seen in studies implementing intrinsically motivated systems [22], [24]. The experiment was repeated twenty times for each object and we performed the Kruskal-Wallis test to analyse the goal distributions. The test was conclusive for the three objects with $H(4),p<0.01$ and the Mann-Whitney U test between two consecutive goals depicts no significant variation in the learning sequence for all of them (see Table I). However, the statistical signification is closer to the baseline experiment than for the inhibition parameters. The persistence parameters provide similar results to baseline for the cube. Regarding the cylinder, the distributions of the different learning progress are more distinct. This manifests less learning overlap between goals across the runs. Regarding the ball, the learning time for all skills is reduced compared to baseline but increased compared to inhibition.
@@
Experimenting on three different objects with various learning complexity allows to evaluate the error inhibition and persistence. Regarding the former, this neural mechanism is introduced to support the switch of goal learning if the learning progress does not evolve. For the latter, we demonstrate that the time window spent on trying to learn a goal can have an influence depending on the object. Together, these parameters demonstrate that they could improve the learning of different skills by relying on the complexity of the objects. Here, the baseline experiment is optimal for the cube with no significant learning fluctuation. The ball exhibits the shortest learning time with moderate learning fluctuation under the inhibition settings and the cylinder displays a weaker overlap of learning between goals under the persistence criterion.

### v2

@@ III-E Curiosity
When the robot discovers a new goal, a forward and an inverse model are created. The forward model outputs the first error to compute the learning progress with dynamic neural fields (Fig. 2). More precisely, the error controller forms a two-dimensional Gaussian input located at the location of the goal angle and object, with an amplitude corresponding to the error value (goal error MT in the Error component). During exploitation, the robot performs a motion and observes the outcome in the environment. In the absence of changes (i.e., the object remains stationary), the error is unchanged. In case of any changes, the error controller computes the error with the forward model. Here, the model follows a precise timing for the error computation, where the observation takes place ahead of the end of the motion. In practice, the robot pushes the object, computes the error while going back to a resting position, and finally signals that the motion is done. The curiosity mechanism is then divided into two parts: the error and learning progress processes. The former focuses on characterizing the error of a goal as neural activity, while the latter computes the error’s level dynamically to deliver the learning progress.
The compute error NF gathers error levels from all the goals. When the robot pursues the learning of a goal, the error level of the goal error MT is combined with the filter MT to affect only the current goal. Indeed, this last field regroups each error and inhibits the current goal, whereas the goal error MT provides a single peak (i.e., the goal error’s activation). It is worth mentioning that the activation function ($\sigma$) used for this field, local goal NF, local error NF, and updated error NF is a ReLU function. The time node acts as a trigger when a new error value is computed. Once a new error is projected by the error controller, this brief activation supports the memory trace update as well as boosting the resting level of other neural fields. The errors MT keeps in memory each goal value and only updates them once the end-of-action node (EOA) is active. The last main fields are the local goal NF and the updated error NF. They have a resting level of 2-2, meaning that they yield peaks of activation only when they receive the boost from the time node. As mentioned earlier, the error is computed before the robot reaches its resting pose and activates the EOA node. This implies that when the time node is active, the local error NF contains the error’s goal before the update, whereas the updated error NF delivers the newly calculated value. These last fields prepare the future neural computation of the various learning progresses.
@@
The learning progress mechanism is straightforward (Fig. 2). It receives inputs from the updated error NF and the local goal NF that are projected onto zero-dimensional nodes with a resting level of 1.025-1.025. The time node remains the same as the error architecture. With this pattern of incoming activation, the rise and low nodes are responsible for detecting if the updated error value is, respectively, superior or inferior to the previous level of activation (local goal NF). The filter NF serves the same purpose as the one described in the error architecture, providing the error level of each goal except the current one. Then, the compute LP memory trace is where the learning progress of a goal is computed. If the low node is active, the suprathreshold activation from the variation NF will project into compute LP and build up an activation. When the rise node is active, it indicates that the error level is rising. In that case, there will be no activation within the variation NF, and the compute LP neural field will decay the learning progress of the current goal. If an activation was present within the variation NF, the activation will increase depending on the value of ($\tau_{+}$). Finally, there is a case where the error values before and after the update are too close to each other, indicating that the progress is not significant. Then, the rise and low nodes will not activate. The specific resting level value of these nodes supports the fine-tuning of this specific event. Here, there will be no activation if the error does not vary beyond a threshold of 0.025.
@@ IV Experimental setup
The purpose of the experiments is to evaluate the learning behavior of the architecture under different parameters and with different objects. More specifically, we want to determine the role of habituation, persistence, and error inhibition during goal discovery and learning. We will also analyze the tonic and phasic activation within the LC. Then, we will assess how the Locus Coeruleus model that gathers bottom-up attention and curiosity leads to an open-ended learning approach with a decentralized cognitive control of goals.
@@
The scene simulates a Franka Emika collaborative robot in Gazebo (Fig. 4). Dynamic neural fields are simulated with the Cedar software [81], and the ROS middleware [82] assures communication between the robot and the neural fields. The objects used for the setup are a red ball, a blue cube, and a yellow cylinder. We specifically choose these objects with regard to their difficulty in manipulating them. Since the end-effector orientation is fixed with the gripper closed, it is relatively easy to push the cube in different directions. For the cylinder, the last pose generated has to be more precise to avoid seeing the object slip away during a motion. The ball is the most difficult object to learn to manipulate, since a slight error in the end effector position can push the ball in various directions. The experiment begins when an object is present in the visual field of the camera. The object is respawned in the center of the table if it is outside the reach of the robot. This means that objects can be moved from various locations on the table as long as they are not out of reach. Each neural field is bounded between 0 and 100 along its respective dimension. The gripper orientation is fixed, as well as the dimension z when performing a motion. During the goal discovery stage, the end effector position of the robot is recorded from the resting state until it reaches the goal pose (III). The number of recorded poses remains almost the same, depending on the motion ($\approx$ 40 poses), and they are used to generate a DMP if the motion leads to new stimuli.
@@
[…]
@@
The learning of a goal is formalized by a forward and inverse model through multilayer perceptrons. The forward model has one input layer (4D), one hidden layer (6D) and one output layer (1D). The inputs are the current position of the object (x,y) and the motor command (final pose of the end effector $x,y$). Then, the output of this neural network is the motion direction (goal angle) of the object. The inverse model has one layer input (3D) composed of the current object position (x,y) and the goal angle of the object. After the hidden layer (4D), the output (2D) is the motor command (end effector position on x and y). Each DMP is associated with its own forward and inverse model. The purpose of the inverse model associated with a DMP is to learn to generate the correct end-effector final pose for a given goal and object position. More details of the learning parameters can be found in Appendix -F.
@@ V-A Habituation
Figure 5: Average number of goals discovered for 10 experiments on each object with fast and slow habituation. The values for fast and slow habituation ($\tau_{+}$) are 2 and 4 seconds.
In this section, we evaluate the effect of habituation during exploration (see Fig. 5). To do so, we performed ten explorations for each of the three objects (cube, cylinder, and ball) by varying the $\tau_{+}$ parameter of the visual memory trace (Eq. 4).
As expected, a slow habituation ($\tau_{+}=4$ seconds) allows the robot to spend more time exploring, leading to an increase in the number of goals discovered (see Fig. 5). We can notice a difference among the objects with the ratio :
@@
Figure 6: Neural activation representing the learning progress evolution in time with standard deviation across 20 experiments for the baseline experiment. The figure depicts the average learning progress for the cube (left), the cylinder (middle), and the ball (right).
@@
Goal sequence Cube Cylinder Ball Cube Cylinder Ball Cube Cylinder Ball
@@
Table I: Mann-Whitney U test between 2 consecutive goals for each object, under each setting and for all experiments.
where $S_{avg}$ and $F_{avg}$ are the average number of goals discovered for a slow and fast habituation. The ratios for the cube, cylinder and the ball are, respectively, 1.53, 2.04 and 1.69. This suggests that, given enough habituation time, the robot can double the number of goals discovered with the cylinder. This can be explained by the orientation of the gripper and the inhibition of return. Since the gripper’s orientation is fixed, only a limited number of actions can trigger the discovery of a new goal. The movement is guided by the inhibition of return around the object’s location, so this inhibition has to be precise and thus requires more trials to generate a meaningful motion. In the following section, we provide an analysis of the learning trajectory of different goals with curiosity.
V-B Goal learning progress
@@ V-B1 Time-course of learning progress between objects
The evolution of learning progress is presented in Figure 6. The experiment is repeated twenty times for each object, and to compare the learning between objects and parameters, we fixed the same four goals. These goals are : push up, push to the left, push to the right, push down. The initialization of the forward model weights determines the first error during goal discovery, which then influences the future goal selection for learning. We consider this experiment as the baseline, with a persistence value (transient action MT node) at $4500ms$ and $100ms$ for $\tau_{+}$ and $\tau_{-}$, respectively. The error inhibition (attempts MT in the persistence module projected on LC:phasic) is settled at $100ms$ for $\tau_{+}$ and $\tau_{-}$, which indicates that there is no error inhibition staying in memory between the selection of a goal. The learning progress (LP MT in the learning progress module) is fixed at $2000ms$ for $\tau_{-}$ and $\tau_{+}$. Regarding the LP, we arbitrarily choose these values because a high $\tau_{-}$ would signify that the goal error takes more time to decay even if the error coming from the forward model is already steady at a low level. We performed a Kruskal-Wallis test to evaluate the significance of goal ordering for each object. For the cube, we take the distribution from $t=[0,1800]s$; for the cylinder, we sample $t=[0,2250]s$; for the ball, the distribution range is $t=[0,6000]s$. The goal ordering analysis aims to demonstrate the influence of error inhibition on learning flexibility. We hypothesize that this specific inhibition (i.e., inspired by [62]) can help disengage from a goal that is too difficult to learn and thus reduce persistence.
The distribution differences between the four goals are significant with $H(4)$, $p<0.01$ among the three objects. Even if the sequence of the four goals is significant among all the runs, we still perform a Mann-Whitney U test (see Tab. I) between two successive goals (push-up $\rightarrow$ push-left, push-left $\rightarrow$ push-right, push-right $\rightarrow$ push-down) to verify a possible overlap between them in certain runs. The statistical test for the cylinder is almost not significant, indicating a different ordering between $up$ and leftleft among several runs. We can conclude that the goals follow the same learning order for the baseline experiment. However, these goals do not reflect a developmental trajectory, as they can be learned independently. Moreover, learning begins with the goal that has the highest error, which depends on the initialization of the synaptic weights of the forward model. The difficulty of a goal to be learned can be evaluated by the time spent until the activation of the learning progress falls below $0.01$. For the ball, it is considerably more difficult to learn the set of goals than for the cylinder and the cube ($\approx$ 6000s). This is essentially due to the precision required to generate a goalgoal (Appendix -B).
@@ V-B2 Winner inhibits similar goals
During exploration with the habituation stage, it is possible for the robot to discover similar goals (see Fig. 8).
Figure 8: (Left) Two similar goals in errors MT (Error mechanism) after goal discovery. (Right) Same two goals in filter MT, but the selected goal for learning is inhibited to compute the error.
For example, the robot can push the cube to the left and push the cube to the left, but slightly shifted by a few degrees upward. In that case, only the goal with the highest error from the forward model will be learned in the future. This comes from the neural dynamics of the filter MT (error module) and, more particularly, from the inhibition of the current goal (see Fig. 8). In fact, the filter MT represents the various discovered goals except the current goal via inhibition of current goal NF. This localized inhibition not only affects the goal peak, but also partially or completely alters any peaks of activation in proximity. This effect depends on the standard deviation ($\sigma$) of the inhibition. A nearby goal will be affected by having its neural activation lowered. This results in a lower neural activation for the similar goal when the current goal sees its error updated. An elevated standard deviation for each neural activation within the neural fields in the Error module would mean that the goals must significantly differ from each other to be learned. The size and amplitude of suprathreshold activations have been shown to have an impact on the learning of sensorimotor contingencies [83]. This effect can be seen as a way to focus first on acquiring a diverse set of goals before learning only slightly different goals later on. The consequences of this effect lead to developmental questions and should be investigated more deeply.
@@ V-C Influence of inhibition and persistence
After learning goals under the baseline parameters, we investigate the impact of error inhibition through the attempts MT in the persistence mechanism. We settle the time parameters at $\tau_{+}=2000ms$ and $\tau_{-}=1500ms$ and run the experiment twenty times for each object (see Fig. 7, top row). We selected these parameters to ensure that the inhibition develops and dissipates more slowly. When a goal error proves challenging to reduce, it induces mild inhibition, preventing the goal from being chosen again in the near future. We repeated a Kruskal-Wallis test for the cube, cylinder, and ball for the time intervals $t=[0,2100]s$, $t=[0,2500]s$ and $t=[0,4250]s$, respectively. The distribution of the four goals in the 20 experiments was conclusive with $H(4),p<0.01$. A Mann-Whitney test was performed to determine whether the goals consistently follow the same sequence (see Table I). We observed that the test is not significant for the cube, which means that there is no clear ordering between pushinguppushing-up and pushingleftpushing-left. If the learning progress for the goal pushinguppushing-up remains stagnant after a learning session, inhibition will hinder near-future selection of this goal. The statistical test result was significant for the cylinder and the ball, but with a higher value than the baseline experiment. This demonstrates the presence of several runs for both objects, where the goals do not follow the mainstream sequence pattern. If we compare the inhibition settings with the baseline, we can see an increase in the learning time for the cube and the cylinder. However, this learning time is shorter for the ball, and the goals pushrightpush-right and pushdownpush-down are less subject to fluctuation. For the ball baseline experiment, pushrightpush-right and pushdownpush-down take, respectively, $5000s$ and $3900s$ against $3000s$ and $2000s$ under the inhibition parameters. The baseline experiment relies only on the learning progress of the goals to continue the exploitation, which provokes only minimal learning fluctuation in the case of easy objects to learn (cube and cylinder). For a difficult object, the system has to continue learning challenging goals exhibiting a certain learning progress with no possibility to select a different goal. Under the inhibition parameters, the model introduces more flexibility in goal selection since learning progress is not the only determining factor. Essentially, this suggests that the robot does not finish learning a goal entirely before learning a new one. This switch between goals indicates an advantage on the ball, whereas it introduces more fluctuation for the cube. Apart from a certain extension of learning time ($\approx 250s$), the cylinder does not appear to be affected by the inhibition parameters.
@@
We introduce the persistence settings by modifying the transient action MT node ($\tau_{+}=6000ms$, $\tau_{-}=100ms$). These settings extend the learning time before the LC module disengages from learning the current goal. This parameter is analogous to the exploration time seen in studies implementing intrinsically motivated systems [24], [26]. Each object was subjected to the experiment twenty times, followed by a Kruskal-Wallis test to evaluate the goal distributions. The test was conclusive for the three objects with $H(4),p<0.01$, and the Mann-Whitney U test between two consecutive goals does not show a significant variation in the learning sequence for all of them (see Table I). However, statistical significance is closer to the baseline experiment than for the inhibition parameters. The persistence parameters provide results similar to the baseline for the cube. With respect to the cylinder, the distributions of the different learning progresses are more distinct. This manifests itself as less learning overlap between the goals in all the runs. With regard to the ball, the learning time for all goals is reduced compared to baseline but increased compared to inhibition.
@@
Investigating three distinct objects of varying learning complexity enables assessing error inhibition and persistence. Error inhibition triggers disengagement from a goal too difficult to learn (i.e., no evolution of the learning progress) and reorients learning to a different goal. Persistence determines the number of trials to learn a goal before disengagement. Together, these parameters modulate the learning of the goals in a flexible way based on the complexity of the objects. Here, the baseline experiment is optimal for the cube with no significant learning fluctuation. The ball exhibits the shortest learning time with moderate learning fluctuation under the inhibition parameters, and the cylinder displays a weaker overlap of learning between goals under the persistence criterion.